# Contact Models in Robotics: a Comparative Analysis

### v2

@@ IV-B Self-consistency of the solvers
Time stepping simulators are sensitive to the choice of the time-step $\Delta t$. Here, we intend to assess the self-consistency of the various contact solvers by examining their deviation when $\Delta t$ grows. Because time discretization also affects the collision detection process, our study is done on the trajectory of a cube dragged on a plane by a growing tangential force and whose contact points should remain constant. This scenario also allows to asses both sticking and sliding modes. For each simulator, a trajectory $\overline{q}$ obtained by simulating the system with a small time-step ($\Delta t=10^{-2}ms$) serves as a reference to compute the state consistency error along the trajectories simulated with larger time-steps (Fig. 15). The deviation from the reference trajectory is quantified via the integral consistency error defined as $\sum_{\tau=0}^{T}\|q^{\tau}-\overline{q}^{\tau}\|\Delta t$.
@@
Stopping criterion. As done in the case of the time-stepping, the sensitivity to the stopping criterion is evaluated by measuring the integral consistency error with respect to a high-resolution ($\epsilon_{abs}=10^{-9}$) reference trajectory (Fig. 15). As demonstrated by Fig. 15, the solvers appear to have different behaviors when their stopping criteria are modified. The gap between trajectories generated via the NCP contact models stays almost constant to a low value. Trajectories obtained via RaiSim significantly deviates when $\epsilon_{abs}$ is increased, while those resulting from CCP grows quickly from a low to a high consistency error. This reveals the importance of setting $\epsilon_{\text{abs}}$ to control the consistency of simulation with CCP. In addition, Fig. 15 demonstrates that the NCP model robustly simulates the mechanical energy evolution, which is not true for the CCP and RaiSim models. In particular, as the accuracy of RaiSim is relaxed, due to its approximation (24), the error on the MDP becomes uncontained leading to a reduced energy dissipation.
Figure 15: Self-consistency w.r.t. stopping criteria when simulating a cube dragged on a plane by a growing tangential force. Dropping the solver accuracy hardly affects the energy evolution when using NCP/PGS. On the contrary, the energy of trajectories from the CCP/ADMM and RaiSim algorithms is more spectacularly impacted.
@@ IV-C Performance benchmarks
Figure 16: Convergence for contact problems on Solo (Top) and Talos (Bottom) robots. PGS is fast to reach a mild accuracy before saturating, while ADMM and second order algorithms can get to higher precision.
@@
Figure 17: Computational timings measured along a trajectory for three robotic systems (c.f. Fig. 8). Solo-12 and Talos are perturbed by applying an external force at their center of mass, while a ball is dropped in the Allegro hand, so the trajectories are not static. The contact solvers are tested in both cold-start (Left) and warm-start (Right) modes. We simulate the same trajectories to evaluate the benefit of warm-starting, but we use the solution of the previous time step as an initial guess. This leads to significant improvements in the computational timings.
@@
As evoked earlier, in addition to being physically accurate, it is also essential for a simulator to be fast, which, in general, constitutes two adversarial requirements. To evaluate the computational footprint of the various solvers, we measure both the number of iterations and the time taken to reach a fixed accuracy on dynamic trajectories involving robotics systems (Fig. 8).
@@
Looking at the number of iterations required to converge (Fig. 16), PGS approaches appear to be reasonably fast to reach mild accuracy ($\epsilon_{\text{abs}}=10^{-5}$) while they eventually saturate before reaching high precision in complex scenarii (Fig. 16, bottom). We show later this can be unsufficient for challenging tasks (Fig. 19). On the other hand, ADMM, Newton and Staggered Projections algorithms can find high accuracy solutions using only a few, but more costly, iterations.
@@
The latter analysis does not account for the per-iteration computational cost, so we report a study on final timings in Fig. 17. When the contact solvers are cold started, we observe that the second-order optimization techniques [32, 35] are less efficient than the PGS solvers and their cheap per-contact iterations (Fig. 17, left). The advanced first-order algorithms like ADMM (Alg. 3) working on the dual CCP problem (20) stands in-between as they leverage the very efficient Featherstone algebra [54] for the computation of the Cholesky factorisation of $G$ (Sec. III-E). However, leveraging the solution from the previous time step to warm-start the solvers — a common strategy in practice — allows for significantly reducing this gap (Fig. 17, right). Therefore, regarding the study of Sec. IV-A, a trade-off appears for algorithms like ADMM, which treat all the contact points globally. In practice, they might be slower than their PGS counterpart while they benefit from better behaviors on ill-conditioned problems.
@@ IV-D MPC for quadruped locomotion
When the terrain is bumpy (roughness of $10^{-1}m$) and slippery ($\mu=0.3$), the locomotion velocity generated from the RaiSim and CCP models significantly deviates from the NCP one (Fig. 19, top right). This can be expected, in light of our previous study, as both the RaiSim and CCP contact models make physical approximations when contact points are sliding (Fig. 19, bottom left). However, we occasionally observe that NCP/PGS also violates the NCP criterion (15) (Fig. 19, bottom left) but for a different reason: PGS was not able to converge before the maximum number of iterations was reached. Therefore, Gauss-Seidel like approaches appear to be sufficient for mild conditions (Fig. 18) but are not robust enough to ensure convergence of the simulation when the locomotion tasks becomes more challenging (Fig. 19). This also causes increased computations from the solvers, particularly for RaiSim (Fig. 19, bottom right). These observations indicate that the low-level choices of the contact solver may induce significant differences in the high-level behaviors of locomotion controllers on complex terrains.

### v3

@@ IV-B Self-consistency of the solvers
Time-stepping simulators are sensitive to the choice of the time-step $\Delta t$. Here, we intend to assess the self-consistency of the various contact solvers by examining their deviation when $\Delta t$ grows. Because time discretization also affects the collision detection process, our study is done on the trajectory of a cube dragged on a plane by a growing tangential force and whose contact points should remain constant (as done in Fig. 11). This scenario also allows to asses both sticking and sliding modes. For each simulator, a trajectory $\overline{q}$ obtained by simulating the system with a small time-step ($\Delta t=10^{-2}ms$) serves as a reference to compute the state consistency error along the trajectories simulated with larger time-steps (Fig. 15).
@@ IV-C Performance benchmarks
Figure 16: Convergence for contact problems on Solo (Top) and Talos (Bottom) robots. The considered contact problems are extracted from one time step of the full trajectory. PGS is fast to reach a mild accuracy before saturating, while ADMM and second-order algorithms can get to higher precision.
@@
Figure 17: Computational timings measured along a trajectory for three robotic systems (c.f. Fig. 8). The represented timings are obtained by averaging on the entire trajectory. The contact solvers are tested in both cold-start (Left) and warm-start (Right) modes. We simulate the same trajectories to evaluate the benefit of warm-starting, but we use the solution of the previous time step as an initial guess. This leads to significant improvements in the computational timings.
@@
As evoked earlier, in addition to being physically accurate, it is also essential for a simulator to be fast, which, in general, constitutes two adversarial requirements. To evaluate the computational footprint of the various solvers, we measure both the number of iterations and the time taken to reach a fixed accuracy on dynamic trajectories involving robotics systems (Fig. 8). This is done on three different robotics scenarios: the quadruped Solo (12-dof) and the humanoid Talos (32-dof) are in a standing position and perturbed by applying an external force (of respectively 10N and 80N) at their center of mass, while a ball is dropped in the Allegro hand (16-dof), so the trajectories are not static.
@@
Looking at the number of iterations required to converge (Fig. 16), PGS approaches appear to be reasonably fast to reach mild accuracy ($\epsilon_{\text{abs}}=10^{-5}$) while they eventually saturate before reaching high precision in complex scenarii (Fig. 16, bottom). We show later this can be insufficient for challenging tasks (Fig. 19). On the other hand, ADMM, Newton and Staggered Projections algorithms can find high-accuracy solutions using only a few, but more costly, iterations. As mentioned earlier, our implementation of Drake’s solver uses a backtracking line search with an Armijo condition while the original algorithm [35] uses a more advanced routine. This difference could lead to degraded performances here which should be kept in mind when interpreting the results.
@@
The latter analysis does not account for the per-iteration computational cost, so we report a study on final timings in Fig. 17. As they work with various contact model hypotheses and thus have different convergence criteria, it is challenging to make a fair comparison between the solvers. Therefore, we choose to run the solvers in the setup they are usually used in practice: the solver is stopped whenever it reaches an absolute convergence threshold ($\epsilon_{\text{abs}}=10^{-6}$) or otherwise, it is early-stopped if a maximum number of iteration ($n_{\text{iter}}=10^{4}$) is reached or stalling is detected via relative convergence criterion ($\epsilon_{\text{abs}}=10^{-8}$). For this reason, Fig. 17 is only informative about the computational cost but not about the accuracy of solvers. When the contact solvers are cold started, we observe that the second-order optimization techniques [32, 35] are less efficient than the PGS solvers and their cheap per-contact iterations (Fig. 17, left). The advanced first-order algorithms like ADMM (Alg. 3) working on the dual CCP problem (20) stands in-between as they leverage the very efficient Featherstone algebra [54] for the computation of the Cholesky factorization of $G$ (Sec. III-E). However, leveraging the solution from the previous time step to warm-start the solvers — a common strategy in practice — allows for significantly reducing this gap (Fig. 17, right). Therefore, regarding the study of Sec. IV-A, a trade-off appears for algorithms like ADMM, which treat all the contact points globally. In practice, they might be slower than their PGS counterpart while they benefit from better behaviors on ill-conditioned problems.
@@ IV-D MPC for quadruped locomotion
When the terrain is bumpy (roughness of $10^{-1}m$) and slippery ($\mu=0.3$), the locomotion velocity generated from the RaiSim and CCP models significantly deviates from the NCP one (Fig. 19, top right). This can be expected, in light of our previous study, as both the RaiSim and CCP contact models make physical approximations when contact points are sliding (Fig. 19, bottom left). However, we occasionally observe that NCP/PGS also violates the NCP criterion (15) (Fig. 19, bottom left) but for a different reason: PGS was not able to converge before the maximum number of iterations was reached. Therefore, Gauss-Seidel-like approaches appear to be sufficient for mild conditions (Fig. 18) but are not robust enough to ensure convergence of the simulation when the locomotion tasks become more challenging (Fig. 19). This also causes increased computations from the solvers, particularly for RaiSim (Fig. 19, bottom right). These observations indicate that the combination of low-level choices on both the contact model and solver may induce significant differences in the high-level behaviors of locomotion controllers on complex terrains.